​if P(-1,0), Q(0,0), R(3,3^(1/2)) are three points then the equation of angle bisector of PQR is

Grade:11

1 Answers

Vikas TU
7 years ago
Slope = tan(theta) b/w the lines PQ and PR would be:
m = tan(theta) = |m1 – m2|/|1 + m1*m2|
where m1 = 0
m2 = (root(3) – 0)/4 = > root(3)/4
Put m1 and m2 values in m to get thetha angle or slope.
thus,
m = tan(theta) = root(3)/4
theta = arctan(root(3)/4)
theta / 2 = [arctan(root(3)/4)]/2
=> 23.4/2 degree
=> 11.7 degree.
new slope for bisector = > tan(11.7) => 0.20 (approx.)
Eqn of bisector,
=> y = (0.20)x